# DLPOLY 5: Calculation of system properties on the fly for very large systems via massive parallelism

### v1

@@ Abstract
keywords:
, Molecular dynamics , Materials modelling , online algorithms , DL_POLYPROGRAM SUMMARY/NEW VERSION PROGRAM SUMMARY
@@
Programming language: Fortran95
@@
Journal reference of previous version: Todorov, I.T., Smith, W., Trachenko, K. and Dove, M.T., 2006. DL_POLY_3: new dimensions in molecular dynamics simulations via massive parallelism. Journal of Materials Chemistry, 16(20), pp.1911-1918.
Does the new version supersede the previous version?: Yes
Reasons for the new version: Development of on-the-fly correlation analysis.
Summary of revisions: General code modularisation, a new control format, SPME re-implementation for per-particle contributions, and a general on-the-fly correlation framework.
@@ 1 Introduction
As for any method, MD simulations have their limitations and associated domain of applicability. Several important ones have been relaxed or lifted through the developments of the last decades; others remain unsolved, including the important issue of small system size.
@@
Running these very large simulations, not only brought scientific insights, but also computational insights. First, the size of a single configuration of 100 million atoms (position data only, in Å with one significant figure) in ASCII is 3.5GB. For comparison, the standard DLPOLY format requires 5 times more data for structural data (15 GB) with an additional 2 times more if velocity data is needed (25GB). In an MD simulation, certain structural and dynamical properties (for example, viscosity or velocity correlation functions) often need over 10,000-100,000 configurations to be computed, requiring usually two orders of magnitude more timesteps in order to get good-quality data. This means that a single trajectory file of this type for a 1 million atom system will require 0.1-1 PB of storage. On the other hand, the work file system of a high-performance facility such as UK’s ARCHER2 service [6] has a total of 10.2 PB [7] to be shared between all users (a maximum of 60 TB allocated to a single project or team). Therefore, the trajectory files of very large systems become impossible to store, let alone process, on high-performance computing facilities.
@@
The problems outlined above are general and will either prevent users from running and analysing MD simulations of large systems or make these simulations inefficient. In our radiation damage case study, we overcame these problems by implementing a special on-the-fly algorithm in the general-purpose MD package, DLPOLY [8]. This algorithm specifically finds and analyses displaced and defect atoms in the very large billion-atom system. This was a non-trivial task as the algorithm had to be made consistent with the domain-decomposition infrastructure of DLPOLY [9, 10]. These simulations have provided a useful case study and a motivation for the development of a more general code, with a view to benefit a wider modelling community and to approach the problems in a general way.
@@
We recall that some properties such as energy can be sampled and averaged at each time step, and these are calculated by DLPOLY and other popular MD codes as a standard output. However, there are other important properties relying on time correlation functions which need to be calculated over the entire trajectory and which are currently calculated by writing out the trajectory file and analysing it a posteriori. One common example is the velocity autocorrelation function related to the phonon density of states and power spectrum of the system. Other examples include k-space densities and currents. For very large system sizes, the problems discussed above make performing these analyses from trajectory files impractical.
Instead of attempting to run a trajectory and subsequently analysing it, we have enabled the calculation of important physical properties on-the-fly.
@@
This sort of on-the-fly calculation will be crucial to the execution and analysis of ever-larger MD simulations. Particularly as there is a growing need to simulate larger systems, length and energy scales, coupled with the availability of scalable MD codes and massive parallel computing facilities.
@@
Supported by the EPSRC funding, we have extended the UK flagship MD code, DLPOLY, to calculate system properties on-the-fly culminating in the DLPOLY 5 code available at [11]. DLPOLY is a general-purpose massively parallel MD simulation package that uses a highly efficient set of methods and algorithms, including domain decomposition, linked cells, Daresbury Advanced Fourier Transform, Trotter derived Velocity Verlet integration and RATTLE [8]. Written to support academic and industrial research, DLPOLY has a wide range of applications and can run on a wide range of computers; from single processor workstations to massively parallel high-performance facilities. During development a strong emphasis was placed on efficient utilization of multi-processor power by optimising memory workload and distribution. This makes it possible to efficiently simulate large systems of billions of atoms. Simulation of systems of this size has been tested, and we are not aware of reasons preventing from running larger systems. DLPOLY 5 is hosted on GitLab [11] has been a free and open source project (GPLv3.0) since 2020.
@@
In this paper, we document and detail the developments of DLPOLY related to on-the-fly calculation of some key system properties. This included general and substantial refactoring across DLPOLY’s modules. We discuss the implementation of on-the-fly calculations of different properties including correlation functions, viscosity and thermal conductivity, elastic constants, currents, and rigid body properties. We then discuss the calculations of these properties and show the results from simulations of large 100-million systems. These results are compared to smaller-size calculations, experiments and previous MD simulations. We begin with an overview of DLPOLY’s history up to DLPOLY 4, including a discussion on its scalability. Then we detail the re-engineering of the code, laying the groundwork for our on-the-fly developments.
2 DLPOLY, massive parallelism and scalability
The DLPOLY project was conceived in 1993 by W Smith and originally released in 1994 as DLPOLY 2 [12, 13] (now known as DLPOLY_CLASSIC [14]) using a replicated data (RD) approach as its underlying parallelisation strategy [15, 16]. Written in Fortran77 [17] and using MPI [18] for interprocessor communication the project brought significant improvements in both functionality and internal algorithms performance in the code, establishing the link-cell (LC) algorithm for efficient search and listing of particle pairs based on a cutoff [19] and the use of 3D discrete fast Fourier transforms (3D FFTs) [20] for the central grid-based operation in the evaluation of the long-ranged contributions of electrostatics interactions [21]. However, by the early 2000s it was clear that RD strategy was rather limited in performance and scalability gains especially for then very large model systems, of the order of 100,000 atoms, and large processor counts, of the order of 100 CPUs, thus necessitating a change in parallelisation strategy and reformulation of its algorithms. The particular deficiencies of the RD approach that needed addressing were the avalanche of communication which increased with the size of a model system (creating larger inter-processor messages) as well as with the processor count (increasing the number of senders and receivers per message for all-to-all communications), and that there was no memory distribution at all, these were major bottlenecks for running large systems of the order of millions of atoms.
In 2003, led by I. Todorov, with core work carried out by W. Smith and I. Bush, DLPOLY 3 was released as part of the fundamental work supported by the eMinerals project [22], brought in as part of the UK’s eScience programme [23]. Not only did this version bring a new code infrastructure in modularised Fortran 90, but it also developed the foundation of the domain decomposition (DD) parallelisation strategy [9, 10], including a Smooth Particle Mesh Ewald (SPME) method suitable for the DLPOLY DD [24], still used in the newer versions. Based on the two premises of: (i) limited density variation - that the model system average particle density does not fluctuate significantly - in space and time, and (ii) the system size/dimensions being much larger than that of the most characteristic interaction cutoff; this strategy developed and addressed the two major shortcomings of the RD approach. First, it provided for near perfect memory distribution as demonstrated in Fig. 2, where three model systems with increasing complexity were run and compared in a weak-scaling manner, i.e. every time the system size doubled, the processor count used to run the simulation doubled. This opened the door to running multi-million atom simulations from its very first release [25].
The weak scaling provides information about the deviation from perfect performance, where it is assumed that doubling processors halves the time-per-atom, i.e. constant speed gain. However, we see in Fig. 2, that even for the most trivial system, Ar, one can see small deviation from the perfect performance due to inter-process communication, necessary for domain halo boundary data exchange. It is clear that, as the model’s complexity increases, so too does the toll on performance, demonstrated by the greater deviation from perfect scaling. For the second model system, NaCl, this is due to long-ranged coulombic interactions, bringing the cost of inclusion of electrostatics into focus. As discussed [24] this is due to communications overheads that are a necessity of the FFT operations needed for evaluation of the long-ranged electrostatics interactions. The third model system, SPC water, is subject to a further performance penalty which is caused by the inclusion of constrained bond dynamics. The reason for this is simple to understand: as more processors are working together, there are more inter-domain faces and hence more constrained oxygen-hydrogen bonds crossing inter-domain faces. Constraining these bonds requires multiple small inter-process updates, which become costly as the model system or processor counts become bigger.
In strong scaling, we track the performance of the key underlying algorithms on the same system size as we increase the processor count. As shown in Fig. 3 the domain decomposition parallelisation strategy provides a good load-balancing and near-perfect (linear) performance, up to to a limit, by the majority of its core algorithms. The speed gain curve deviates from its linear form only for the long-range interactions (Ewald k-space). This evaluation is also the ultimate cause for the cumulative overall speed gain to deviate in a non-linear manner as the processor count increases and decay with increasing the processor count. More recent strong-scaling benchmarking on Archer 2 [6] demonstrated the same excellent scalability for large Ar and NaCl systems. A comprehensive survey of DLPOLY performance on modern hardware can be found in [26].
In 2010 DLPOLY_4 was released, including a major rewrite providing an extension of particle dynamics to rigid body dynamics [27], inclusion of two-temperature thermostat functionality [28] for high-energy events and the inclusion of parallel I/O [29]. The rigid body dynamics provided a natural extension of the parallel performance capabilities for biological and biochemical systems in which constrained bond dynamics was a bottleneck (as seen in the fast decaying weak performance of the SPC water system in Fig. 2). The enhancement in performance of rigid body (RB) dynamics and constrained bond (CB) dynamics is demonstrated in Fig. 4 where both weak and strong scaling tests were performed on the same SPC water system and model with scaled sizes in the case of weak scaling. The difference between RB and CB dynamics is the way in which bond constraints are applied. In the CB case, it is based on simple, but iterative, application of small constant force updates on top of the Newtonian equations of motion, which in DD requires frequent local inter-process data exchanges. In contrast, in the RB case the Eulerian equations of RB rotation are solved alongside with the Newtonian ones. However, these further calculations, despite their complexity, are computationally cheaper (no iterative computation) and require no iterative communication for a difference with CB dynamics one. Further developments in 2012 led to the inclusion of Dissipative Particle Dynamics integration [30] and a three seeded random number generation [31] for using repeatable random series for stochastically dependent thermostating, kinetic fields and pressure tensors in a fully deterministic manner across domains and time series.
2.1 Current state
A few years ago an extensive refactoring effort took place that is the the base of current DLPOLY. These changes were introduced mainly due to the need for being able to add new complex tasks with maximum code reuse and allow various easy optimisations. The main changes introduced were: i) using object oriented Fortran – 2018 as the minimal standard required, ii) restructuring the code to be fully modular, with all functions and subroutines exchanging data only via parameters rather than via modules - removal of implicit common blocks. Only one singleton remains for dealing with I/O for errors. iii) introduced an automated test system with both regression and unit tests. Testing was extended from 24 tests to over 200 tests. iv) atomic configurations are stored in an array of structures that is cache friendly with respect to positions and forces, v) electrostatics was rewritten to allow per-particle decomposition and allowing the computation of long range contributions from any power potential, see next section for details. iv) The CONTROL file format was redesigned (see appendix C). In general all entries now are consistently of the form token-value-unit (where applicable), STATIS, RDF and other output files can be written in the yaml format [32] for easier postprocessing, v) a modern software infrastructure is in place with full CI/CD via gitlab.com, vi) license was changed to an open-source license GPL-3.0, vii) dlpoly-py was developed [33] - a companion python package that can read inputs and outputs of DLPOLY and initiate simulations. Without this extensive refactoring, implementation of the Empirical Valence Bond [34] and the new Chemshell interface [35] would not have been possible.
2.2 SPME refactoring
As part of the DLPOLY5 developments, we have recently benchmarked 2-body and SPME-enabled systems composed of million atoms and 2-body simulations composed of billion atoms across increasing MPI process counts on Archer2 [6]. We examine processor counts ranging from 128 in the small systems to 262,144 in the larger system. Representing a 35% concurrent utilisation of Archer2’s 750,080 CPU cores [38]. The results can be found in Fig. 5. Primarily the results continue to demonstrate the scalability of DLPOLY. But secondly, it also demonstrated the scalability of the newly refactored MD algorithms, with the reworked DD memory distribution that favoured purely local communications and minimised all-to-all communications as much as possible. A comprehensive survey of DLPOLY performance on modern hardware can be found in [26].
@@ 3.1 On-the-fly methods
In MD, the configuration of atomic positions is mapped through simulation time-steps into a trajectory. In order to calculate any properties defined by these time-linked configurations the trajectory may be stored in memory, and an algorithm applied. For example storing atom velocities and calculating the velocity auto-correlation function (VAF). For small systems, it may be possible to store these data at every step, obtaining a VAF accurate to the time-step . As the system size scales this becomes infeasible both in terms of storage and I/O penalties if storing to disc. By conceptualising the simulation trajectory as a big-data stream, we can apply the same ideas of online algorithms to calculate system properties on-the-fly. For modest systems, this presents a convenience. However for large systems, this methodology can give access to properties which otherwise require infeasible storage requirements or some accuracy trade-off.
@@ 3.2 Correlation functions at runtime
Given two observable quantities and in MD, where are simulation time steps for a time step of , we can evaluate the correlation function of these two values at time-lags as [43],
@@
This approach is similar to the velocity correlator introduced by Frenkel and Smit [2] with the additional separation of (data points stored at each level) and the averaging parameter. By increasing and (with fixed), later correlation times may be collected at a fixed accuracy. Increasing has the effect of calculating much longer correlation times for a trade off in accuracy and increased performance. The maximum lag-time of a correlation given these parameters is defined by where is the simulation time step and the frequency the correlation is updated is . The trade off in performance terms scales as . The actual performance will also depend on the values correlated (e.g. velocity across all atoms, or system stress) and other system options (such as electrostatics).
@@
During the simulation, correlators are stored index-able by the unique names of the observables they correlate, i.e. at any point in the simulation, we may trivially obtain the current correlation value for say the heat-flux in , heat_flux_x-heat_flux_x. Not only does this serve to maintain uniqueness, but also allows us to easily report these values, or process them further, at runtime. The result of each correlator is reported in a new file COR which reports the value of each correlation function as well as potentially a set of derived quantities, at user specified intervals. These derived quantities include viscosity, thermal conductivity, and elastic constants as detailed in the next sections.
@@ 3.3.1 Comparison to experiment
For all simulation work, we began by equilibrating for steps in an NPT ensemble, followed by steps with an NVE ensemble to account for changing thermostats, and then production runs, where statistics are calculated, ran for time steps. In all cases the time step was ps. For correlations we collect data for lag times up to ps. For the small systems () we collect data on independently seeded simulations for replicates to ensure adequate averaging of the Green-Kubo results. For the large-scale systems, we were able to simulate replicates for each data point.
@@ 3.4.1 Comparison to simulations and experiments
After equilibration production runs (NVT) were simulated for steps unless otherwise specified. For these runs, the atoms where re-scaled to the average simulation cell from the equilibration phase. Statistical averages were calculated on a rolling window of ps updated every step, and correlations computed for a maximum lag time of ps, also updated every step using DLPOLY’s on-the-fly correlators. Everything else was kept the same. For each replicate , this procedure was carried out with a different random seed for equilibration and production. The methodology is similar to [61], except they generate initial inputs through Monte-Carlo (MC) methods.
@@ 3.6 Rigid body correlations
Using this on-the-fly correlation functionality we are able to reproduce the results of Brodka and Zerda [69] for SF6. We do this by simulating rigid bodies of SF6 molecules with the same Lennard-Jones potential parameters. For equilibration we set the temperature to 296 K, and pressure to the experimental values listed for the pressure by Brodka and Zerda. We simulate in NPT for ps to ensure equilibration. The resulting densities were similar to the and g/cm3 listed in their results. We then rescaled our resulting configurations to match those densities exactly, and then proceed in NVE collecting statistics over ns simulation steps. Brodka and Zerda use equilibration steps and production steps at a higher time step ( fs vs. fs). Our results for the VAF are shown in Figure 11 which match well.

### v2

@@ Abstract
PROGRAM SUMMARY/NEW VERSION PROGRAM SUMMARY
@@
CPC Library link to program files: (to be added by Technical Editor)
@@
Programming language: Fortran 2008
@@ 1 Introduction
As for any method, MD simulations have their limitations and associated domain of applicability. Several important ones have been relaxed or lifted through the developments of the last decades; others remain unsolved, including the impact of small system size on on simulated properties, such as diffusion coefficients [4].
@@
Running these very large simulations, not only brought scientific insights, but also computational insights highlighting the scale of the data. In single precision, the uncompressed positions of 100 million atoms for one timestep costs approximately 1.2GB of disc storage. Computed velocity and force data brings this to 3.6GB. In an MD simulation, certain structural and dynamical properties (for example, viscosity or velocity correlation functions, momentum currents, and radial distribution functions) often need over 10,000-100,000 configurations to be computed, requiring usually two orders of magnitude more timesteps in order to get good-quality data. For such analysis the data requirement reaches 100’s of Terabytes. On the other hand, the work file system of a high-performance facility such as UK’s ARCHER2 service [7] has a total of 10.2 PB [8] to be shared between all users (a maximum of 60 TB allocated to a single project or team). Going beyond to 1 billion atoms will require Petabytes of data and yet simulations of this magnitude are increasingly being performed [9, 10, 11, 12]. The magnitude of these storage requirements presents a significant storage challenge which must be addressed before any post-processing to obtain results.
@@
In addition to these problems many MD use-cases rely on the analysis of rare events such as protein folding or phase transitions. Both accessing these events and the analysis of them in MD exacerbates the issue of trajectory storage. In this domain the exploitation of collective variables (CV) [13] both to inform sampling or otherwise bias dynamics and to detect or drive these events is well used. As a result software such as PLUMED [14, 15, 16] and A4MD [17] act as in situ monitors of MD simulations to ingest data trajectory data at runtime to monitor CVs and act upon them.
The problems however extend more generally and will either prevent most users from running and analysing MD simulations of large systems or make these simulations inefficient. In general this challenge can be partially addressed with multiple methodologies related to data compression. Plain ASCII is often a default and the simplest schemes merely reduce written significant figures (for example this is optional in DLPOLY, and also with GROMACS’s XTC format [18]). Nevertheless, this comes at potentially significant accuracy losses. Alternatively common lossless formats such as GZIP are also offered, for example in LAMMPS [19] which can achieve significant file size reductions. However more advanced algorithms specifically for floating-point data exists such as ZFP [20]. This compression can be performed at a relatively low cost including into efficient domain specific data-layouts such as H5MD [21] that utilises HDF5 [22] for MD data. For example lossless compression ratios of around 1.44 with ZFP compared with 1.09 for GZIP on the same test data suite can be achieved. With lossy compression this rises up to 10.98 or 226.3 with 32-bit or 16-bit precision. ZFPs throughput rates of around 20 MB/s to 60 MB/s for 64 and 32 bit precision on single CPU cores makes this a negligible cost [20]. In the lossless case this can be expected to result in a 30% reduction in I/O overheads. If 32-bit (or less) lossy precision is acceptable using ZFP compressed trajectory output and on-the-fly methods will likely perform similarly in terms of workflow core hours. Since the data foot-print will be reduce by around 90% (or more). Note however that the on-the-fly method involves no precision loss by bit-per-float reduction or lossy compression. It may also be used in conjunction with a more manageable lossy trajectory write-out. For extremely large systems lossless storage and post-processing analysis are still problematic.
@@
Alternatively these problems can be addressed by bypassing trajectory storage entirely. In the simplest cases observables can trivially be computed each step, on-the-fly, without (or whilst) saving the trajectory such as the stress tensor. A more complex example is our radiation damage case study - we overcame these problems by implementing a special on-the-fly algorithm in the general-purpose MD package, DLPOLY [23, 24]. This algorithm specifically finds and analyses displaced and defect atoms in the very large billion-atom system. This was a non-trivial task as the algorithm had to be made consistent with the domain-decomposition infrastructure of DLPOLY [25, 26]. These simulations have provided a useful case study and a motivation for the development of a more general code, with a view to benefit a wider modelling community and to approach the problems in a general way.
@@
Recall that some properties such as energy can be sampled and averaged at each time-step, and these are calculated by DLPOLY and other popular MD codes as a standard output. However, there are other important properties relying on time correlation functions which need to be calculated over the entire trajectory and which are currently calculated by writing out the trajectory file and analysing it a posteriori. One common example is the velocity autocorrelation function related to the phonon density of states and power spectrum of the system. Other examples include k-space densities and currents[27]. For very large system sizes, the problems discussed above make performing these analyses from trajectory files impractical.
@@
Instead of attempting to save a trajectory and subsequently analysing it, we have enabled the calculation of important physical properties on-the-fly. This sort of on-the-fly calculation will be crucial to the execution and analysis of ever-larger MD simulations, particularly as there is a growing need to simulate larger systems, length and energy scales, coupled with the availability of scalable MD codes and massive parallel computing facilities.
@@
Supported by the EPSRC funding, we have extended the UK flagship MD code, DLPOLY, to calculate system properties on-the-fly culminating in the DLPOLY 5 code available at [28]. DLPOLY is a general-purpose massively parallel MD simulation package that uses a highly efficient set of methods and algorithms, including domain decomposition, linked cells, Daresbury Advanced Fourier Transform, Trotter derived Velocity Verlet integration and RATTLE [24]. Written to support academic and industrial research, DLPOLY has a wide range of applications and can run on a wide range of computers: from single processor workstations to massively parallel high-performance facilities. During development a strong emphasis was placed on efficient utilization of multi-processor power by optimising memory workload and distribution. This makes it possible to efficiently simulate large systems of billions of atoms. Simulation of systems of this size has been tested, and we are not aware of reasons preventing from running larger systems. DLPOLY 5 is hosted on GitLab [28] has been a free and open source project (GPLv3.0) since 2020.
@@
In this paper, we document and detail the developments of DLPOLY related to on-the-fly calculation of some key system properties. This included general and substantial refactoring across DLPOLY’s modules to enable the new developments. We begin by discussing the theory and implementation of on-the-fly calculations of correlation functions. We then show results from simulations of small and large scales - up to 100 million atoms - for calculating different properties including viscosity and thermal conductivity, elastic constants, currents, and rigid body properties. These results are compared with experimental results and previous MD simulations. After showing the case studies we discuss benchmarks of our work alongside the costs and benefits or using on-the-fly methods in DLPOLY, and more generally. We finish with an overview of DLPOLY’s recent history since our last published work [24], the 4th and now 5th major releases. Including in particular the refactoring which laid the groundwork for the on-the-fly developments.
@@ 2.1 On-the-fly methods
In MD, the phase space configuration of atoms is mapped through simulation time-steps into a trajectory. In order to calculate any properties defined by these time-linked configurations the trajectory may be stored on disc, and an algorithm applied. For example storing atom velocities and calculating the velocity auto-correlation function (VAF). For small systems, it may be possible to store these data at every step, obtaining a VAF accurate to the time-step . As the system size scales this becomes infeasible both in terms of storage and I/O penalties if storing to disc. By conceptualising the simulation trajectory as a big-data stream, we can apply the same ideas of online algorithms to calculate system properties on-the-fly. For modest systems, this presents a convenience. However for large systems, this methodology can give access to properties which otherwise require infeasible storage requirements or some accuracy trade-off.
@@ 2.2 Correlation functions at runtime
Given two observable quantities and in MD, where are simulation time-steps for a time-step of , we can evaluate the correlation function of these two values at time-lags as [33],
@@
This approach is similar to the velocity correlator introduced by Frenkel and Smit [2] with the additional separation of (data points stored at each level) and the averaging parameter. By increasing and (with fixed), later correlation times may be collected at a fixed accuracy. Increasing has the effect of calculating much longer correlation times for a trade off in accuracy and increased performance. The maximum lag-time of a correlation given these parameters is defined by where is the simulation time-step and the frequency the correlation is updated is . The trade off in performance terms scales as . The actual performance will also depend on the values correlated (e.g. velocity across all atoms, or system stress) and other system options (such as electrostatics).
@@
During the simulation, correlators are stored indexed by the unique names of the observables they correlate, i.e. at any point in the simulation, we may trivially obtain the current correlation value for say the heat-flux in , heat_flux_x-heat_flux_x. Not only does this serve to maintain uniqueness, but also allows us to easily report these values, or process them further, at runtime. The result of each correlator is reported in a new file COR which reports the value of each correlation function as well as potentially a set of derived quantities, at user specified intervals. These derived quantities include viscosity, thermal conductivity, and elastic constants as detailed in the next sections.
@@ 2.3.1 Comparison to experiment
For all simulation work, we began by equilibrating for steps in an NPT ensemble, followed by steps with an NVE ensemble to account for changing thermostats, and then production runs, where statistics are calculated, ran for time-steps. In all cases the time-step was ps. For correlations we collect data for lag times up to ps. For the small systems () we collect data on independently seeded simulations for replicates to ensure adequate averaging of the Green-Kubo results. For the large-scale systems, we were able to simulate replicates for each data point.
@@ 2.4.1 Comparison to simulations and experiments
After equilibration production runs (NVT) were simulated for steps unless otherwise specified. For these runs, the atoms were re-scaled to the average simulation cell from the equilibration phase. Statistical averages were calculated on a rolling window of ps updated every step, and correlations computed for a maximum lag time of ps, also updated every step using DLPOLY’s on-the-fly correlators. Everything else was kept the same. For each replicate , this procedure was carried out with a different random seed for equilibration and production. The methodology is similar to [50], except they generate initial inputs through Monte-Carlo (MC) methods.
@@ 2.6 Rigid body correlations
Using this on-the-fly correlation functionality we are able to reproduce the results of Brodka and Zerda [58] for SF6. We do this by simulating rigid bodies of SF6 molecules with the same Lennard-Jones potential parameters. For equilibration we set the temperature to 296 K, and pressure to the experimental values listed for the pressure by Brodka and Zerda. We simulate in NPT for ps to ensure equilibration. The resulting densities were similar to the and g/cm3 listed in their results. We then rescaled our resulting configurations to match those densities exactly, and then proceed in NVE collecting statistics over ns simulation steps. Brodka and Zerda use equilibration steps and production steps at a higher time-step ( fs vs. fs). Our results for the VAF are shown in Figure 7 which match well.